OpenStudy (anonymous):

Simplify. –15 – (–3) + 4

OpenStudy (anonymous):

-15-(-3)=-15+3=12 So –15 – (–3) + 4 =?

OpenStudy (anonymous):

I mean -12

OpenStudy (anonymous):

-12--3+4=-5

OpenStudy (anonymous):

-15--3+4=-8

OpenStudy (anonymous):

Now award metal

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

now this one Simplify. 5 • 2 • (–4) + 3 • 6